# Optimal Projection for 3D Gaussian Splatting

### v2

@@ Abstract
3D Gaussian Splatting has garnered extensive attention and application in real-time neural rendering. Concurrently, concerns have been raised about the limitations of this technology in aspects such as point cloud storage, performance , and robustness in sparse viewpoints , leading to various improvements. However, there has been a notable lack of attention to the projection errors introduced by the local affine approximation inherent in the splatting itself, and the consequential impact of these errors on the quality of photo-realistic rendering. This paper addresses the projection error function of 3D Gaussian Splatting, commencing with the residual error from the first-order Taylor expansion of the projection function $\phi$. The analysis establishes a correlation between the error and the Gaussian mean position. Subsequently, leveraging function optimization theory, this paper analyzes the function’s minima to provide an optimal projection strategy for Gaussian Splatting referred to Optimal Gaussian Splatting. Experimental validation further confirms that this projection methodology reduces artifacts, resulting in a more convincingly realistic rendering.
@@
3D deep learning, view synthesis, radiance fields, 3D gaussians, real-time rendering, error analysis
Figure 1: By minimizing the projection error through error analysis, we have achieved an improvement in the rendering image quality compared to the original 3D-GS [16].
@@ 1 Introduction
Novel view synthesis aims to generate new views of a 3D scene or object by interpolating from a set of images with known camera parameters, which is a longstanding challenge in computer vision and a prerequisite for many augmented reality (AR) and virtual reality (VR) applications. From traditional techniques such as Structure-From-Motion (SFM) [31] or Image-Based Rendering (IBR), to neural network-based scene representations [15, 30, 8, 24, 29, 32], various approaches have struggled to address the problem of synthesizing a new view from captured images. Significant progress in this field emerged with the advent of Neural Radiance Fields [25] based on MLPs (Multilayer Perceptrons). Subsequent works have extensively utilized neural radiance fields, including MipNeRF [1] to address aliasing and over blurriness in Neural Radiance Fields, and DietNeRF [14], RegNeRF [28], FreeNeRF [40], etc., to enhance the robustness of NeRF in sparse-view scenarios. NeRF++ [42] analyzes the success factors of NeRF and introduces the concept of Foreground-Background NeRF. Barron et al. introduced MipNeRF360 [2] as an extension of MipNeRF [1] to address the issue of generating low-quality renderings for unbounded scenes in NeRF. However, these methods rely on MLPs with a relatively large number of layers, leading to lengthy training and rendering times, which still fall significantly short of real-time rendering. As a result, many methods resort to auxiliary data structures for acceleration [26, 37, 7, 36, 39]. However, due to the implicit representation of scenes and the dense sampling of points along rays, they still face challenges in achieving real-time performance.
Recently, 3D Gaussian Splatting (3D-GS) [16] has departed from the implicit scene representation using MLPs and instead opted for an explicit representation using Gaussian functions. This approach avoids the need for sampling points during ray marching for rendering integration, replacing it with finite Gaussian splatting into the image space and achieving real-time rendering. Subsequently, a cascade of works and applications has emerged building upon the foundation of 3D-GS. SuGaR [11] is used for efficient 3D mesh reconstruction, aligning surfaces by regularizing geometry through the alignment of Gaussians. Recently, there have been efforts to enhance the robustness in sparse-view scenarios [38], performance, and storage efficiency [27] of 3D Gaussian Splatting. However, these improvement efforts have not specifically addressed the errors associated with Gaussian projection itself. A potential method that involves improvements to Gaussian functions is [41]. However, it introduces a Gaussian filter without investigating the projection function and projection errors. The approach still relies on projecting onto the $z=1$ plane.
Similar to the NeRF [25], the input to 3D Gaussian Splatting (3D-GS) consists of a set of images, together with the corresponding cameras calibrated by SfM [31]. However, in contrast to NeRF, 3D Gaussian Splatting takes the sparse point cloud generated during the SfM calibration process as input. From these points, it constructs a set of 3D Gaussians as graphical primitives explicitly representing the scene. Each 3D Gaussian is characterized by its position (mean) $\boldsymbol{\mu}$, covariance matrix $\boldsymbol{\Sigma}$, and opacity $\alpha$, and carries anisotropic spherical harmonics (SH) representing directional appearance component (color) of the radiance field, following standard practice [26, 7]. Subsequently, these 3D Gaussians are projected onto the image plane ($z=1$ plane) via a projection function denoted as $\phi$ for differentiable rasterization. This process, devoid of NeRF’s ray marching and points sampling for rendering integral, leads to a significant performance improvement compared to the NeRF based methods, achieving real-time rendering.
@@
Unfortunately, Gaussian functions, despite retaining their Gaussian properties through convolution or affine transformation, may not necessarily maintain such characteristics after undergoing a projection transformation. Therefore, 3D Gaussian Splatting adopts a local affine approximation [44], specifically approximating the projection function with the first two terms of its Taylor expansion. Nevertheless, approximations introduce corresponding errors, and the local affine approximation similarly contributes to these errors which may lead to artifacts in the rendered images. This paper exploits the relationship between the error of 3D Gaussian Splatting and the Gaussian mean through the analysis of the Taylor remainder term. Additionally, it identifies the circumstances under which the error is minimized by determining the extremum of the error function.
Finally, based on the extremum analysis of the error function, we propose an optimal projection method. Specifically, we project along the direction from the Gaussian mean to the camera center, where the projection plane is tangent to the line connecting the Gaussian mean and the camera center. Further validation has been conducted on diverse datasets, comparing the proposed projection with the straightforward approach of projecting 3D Gaussians onto the $z=1$ plane. The results illustrate a reduced expected error, leading to real-time, realistic rendering, as shown in Figure 5 and Table 1.
To summarize, we provide the following contributions:
@@
  • We analyze error of 3D Gaussian Splatting that may lead to artifacts and consequently degrades rendering quality. We have identified the correlation between this error and the Gaussian position.
@@
  • Based on the conclusion regarding the extremum of the expectation of this error function, we propose a new optimal projection method. Instead of naively projecting all Gaussians onto the same plane, we now adopt different tangent plane projections based on the directions from each Gaussian mean to the camera center. Through experiments, we further validated the superiority of this method. In various datasets and scenes, it consistently achieves better overall performance compared to the original projection method.
@@ 2 Related Work
To begin with, we will provide a succinct overview of traditional approaches in novel view synthesis and scene reconstruction. Subsequently, we will delve into works related to implicit scene representation using Neural Radiance Fields and explicit scene representation through 3D Gaussian Splatting.
2.1 Traditional Scene Reconstruction and Rendering
Early novel-view synthesis methods were based on light fields  [10, 19], transitioning from densely sampled to unstructured capture  [3]. Structure-from-Motion (SfM) [33] introduced a new era using photo collections for view synthesis. Multi-view stereo (MVS) and subsequent algorithms  [9, 5] blended input images but faced challenges. Recent neural rendering algorithms [18] have overcome these issues, outperforming traditional methods.
@@ 2.2 Neural Radiance Fields
The Neural Radiance Field (NeRF) [25] stands out as a successful neural rendering method based on MLPs, primarily owing to its encoding of position and direction. This encoding allows for effective reconstruction of high-frequency information in scenes. Notable improvements on this encoding have been made by MipNeRF [1], NeRF-W [22], FreeNeRF [40], and Instant NGP [26]. These enhancements enable the handling of multi-resolution image inputs, multi-illumination with occlusion image inputs, sparse-view inputs, and achieve nearly real-time rendering capabilities, respectively. Importantly, these methods thoroughly exploit the intrinsic capabilities of NeRF as an implicit scene representation without introducing additional model priors. However, these methods are constrained by the dense sampling of rays, preventing them from achieving real-time performance. In contrast, the recent 3D Gaussian Splatting [16] achieves real-time rendering by breaking away from MLPs and ray sampling constraints.
@@
3D Gaussian Splatting [16] utilizes Structure-from-Motion (SFM) [31] point clouds as input to construct an explicit scene representation using Gaussian functions. Subsequently, these Gaussian functions undergo a series of coordinate transformations and are projected onto the image plane at $z=1$, where rasterization takes place.
The significant breakthrough in performance is achieved by eschewing MLPs and ray sampling in favor of Gaussian functions and Gaussian Splatting. This shift has garnered widespread attention in the industry, leading to various works building upon this method. Gaussian Splatting is applied in SLAM (Simultaneous Localization and Mapping) systems [23]. Additionally, there is scene editing [6] and segmentation [4] based on 3D Gaussian.
There are also efforts to integrate 3D Gaussian Splatting with the popular diffusion models [13, 34, 35] in the AIGC domain [20]. The diffusion model is employed to provide geometric priors for 3D Gaussian Splatting, correcting inconsistencies in multi-view geometry. Another method involves tracking dynamic Gaussians through continuous dynamic view synthesis [21]. However, these enhancement efforts have relatively less emphasis on the 3D Gaussian Splatting method itself and often involve the use of additional models or priors to assist 3D-GS.
This paper aims to explore the potential of 3D Gaussian Splatting by analyzing the errors that may arise during the projection process. The analysis will delve into the factors contributing to these errors and propose methods to mitigate and reduce them.
@@ 3 Preliminaries
(a) 0.095
(b) 0.95
Figure 2: Visualizations of the 3D Gaussian Splatting error function $\epsilon\left(\theta_{\mu},\phi_{\mu}\right)$. The two subplots from left to right respectively depict the function graphs within the domains of $\left\{\left(\theta_{\mu},\phi_{\mu}\right)\mid\theta\in\left[-\pi/4\cdot 0.095,\pi/4\cdot 0.095\right]~\land~\phi\in\left[-\pi/4\cdot 0.095,\pi/4\cdot 0.095\right]\right\}$ and $\left\{\left(\theta_{\mu},\phi_{\mu}\right)\mid\theta\in\left[-\pi/4\cdot 0.95,\pi/4\cdot 0.95\right]~\land~\phi\in\left[-\pi/4\cdot 0.95,\pi/4\cdot 0.95\right]\right\}$ where $\pi/4$ corresponds to the integration region of the mathematical expectation. These representations are employed for a better understanding of the nature of this error function.
In the world coordinate system, let there be a 3D Gaussian function $G$, characterized by its mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$, expressed as:
$G_{\mathbf{\mu},\boldsymbol{\Sigma}}\left(\mathbf{x}\right)=\exp\left(-1/2\left(\mathbf{x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left(\mathbf{x}-\boldsymbol{\mu}\right)\right)\text{.}$ (1)
This Gaussian is multiplied by $\alpha$ in blending process.
@@
To project the Gaussian onto the image plane, the initial step in 3D Gaussian Splatting involves an affine transformation of this Gaussian function from the world coordinate system to the camera coordinate system via a viewing transformation matrix $\mathbf{W}$. The transformed Gaussian function is as follows:
@@
$\begin{split}G\left(\mathbf{x}\right)&=\exp\left(-1/2\left(\mathbf{x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left(\mathbf{x}-\boldsymbol{\mu}\right)\right)\\ &=\exp\left(-1/2\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)^{\top}\left(\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}\right)^{-1}\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)\right)\end{split}$ (2)
which can also be represented by the following function:
$\begin{split}G^{{}^{\prime}}\left(\mathbf{x^{{}^{\prime}}}\right)&=\exp\left(-1/2\left(\mathbf{x^{{}^{\prime}}}-\boldsymbol{\mu^{{}^{\prime}}}\right)^{\top}{\boldsymbol{\Sigma}^{{}^{\prime}}}^{-1}\left(\mathbf{x^{{}^{\prime}}}-\boldsymbol{\mu^{{}^{\prime}}}\right)\right)\end{split}$ (3)
@@
where $G^{{}^{\prime}}$, $\mathbf{x}^{{}^{\prime}}=\mathbf{W}\mathbf{x}$, $\boldsymbol{\mu}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}$ represent the 3D Gaussian function, point, mean and covariance matrix in the camera coordinate system, respectively.
Subsequently, it is necessary to project the Gaussian function in the camera coordinate system onto the $z=1$ plane. The projection plane is formulated as:
$\mathbf{x_{0}}^{\top}\cdot\left(\mathbf{x}^{{}^{\prime}}-\mathbf{x_{0}}\right)=0$ (4)
@@
where $\mathbf{x_{0}}=\left[\begin{matrix}0,0,1\end{matrix}\right]^{\top}$ represents the projection of the camera coordinate system origin onto this plane. According to the projection plane equation (Equation 4), the projection function $\phi$ is obtained as:
$\phi\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x_{0}}^{T}\mathbf{x_{0}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}.$ (5)
Expanding this projection function to the first order using Taylor series yields:
@@
$\begin{split}\phi\left(\mathbf{x}^{{}^{\prime}}\right)&=\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)+R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\\ &\approx\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (6)
@@
where $\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ is the Jacobian of the affine approximation of the projective transformation, denoted as $\mathbf{J}$. Applying this local affine approximation by neglecting the Taylor remainder term allows to derive the 2D Gaussian function $G_{\text{2D}}$ projected onto the $z=1$ plane similar to Equation 2:
@@
$\begin{split}G_{\text{2D}}\left(\mathbf{x}^{{}^{\prime}}\right)&=\exp\left(-1/2\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\\ &\approx\exp\left(-1/2\left(\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right).\end{split}$ (7)
Note that since the rank of the matrix $\mathbf{J}$ is 2, the inverse of $\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ is, in fact, the inverse of the covariance matrix of the 2D Gaussian which means skipping the third row and column of this matrix. Similarly, this function can be expressed in another form:
$G_{\text{2D}}\left(\mathbf{x_{\text{2D}}}\right)=\exp\left(-1/2\left(\mathbf{x_{\text{2D}}}-\boldsymbol{\mu_{\text{2D}}}\right)^{\top}\boldsymbol{\Sigma}_{\text{2D}}^{-1}\left(\mathbf{x_{\text{2D}}}-\boldsymbol{\mu_{\text{2D}}}\right)\right)$ (8)
where $G_{\text{2D}}$, $\mathbf{x}_{\text{2D}}=\phi\left(\mathbf{x}^{{}^{\prime}}\right)$, $\boldsymbol{\mu}_{\text{2D}}=\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ and $\boldsymbol{\Sigma}_{\text{2D}}=\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ represent the 2D Gaussian function, point, mean and $\times 2$ 2 covariance matrix in the image coordinate system, respectively.
After the aforementioned steps, 3D Gaussian Splatting projects the 3D Gaussian from the world coordinate system onto the image plane. Subsequently, rasterization is performed on the image plane to obtain the rendered image.

### v4

# https://letianhuang.github.io/op43dgs/

@@ Abstract
3D Gaussian Splatting has garnered extensive attention and application in real-time neural rendering. Concurrently, concerns have been raised about the limitations of this technology in aspects such as point cloud storage, performance, and robustness in sparse viewpoints, leading to various improvements. However, there has been a notable lack of attention to the fundamental problem of projection errors introduced by the local affine approximation inherent in the splatting itself, and the consequential impact of these errors on the quality of photo-realistic rendering. This paper addresses the projection error function of 3D Gaussian Splatting, commencing with the residual error from the first-order Taylor expansion of the projection function. The analysis establishes a correlation between the error and the Gaussian mean position. Subsequently, leveraging function optimization theory, this paper analyzes the function’s minima to provide an optimal projection strategy for Gaussian Splatting referred to Optimal Gaussian Splatting, which can accommodate a variety of camera models. Experimental validation further confirms that this projection methodology reduces artifacts, resulting in a more convincingly realistic rendering.
@@
3D deep learning view synthesis radiance fields 3D gaussians real-time rendering error analysis
Figure 1: By minimizing the projection error through error analysis, we have achieved an improvement in the rendering image quality compared to the original 3D-GS [17]. These images are cropped from the complete images of the Train scene, in order to be highlighted.
@@ 1 Introduction
Reconstructing 3D scenes from 2D images and synthesizing novel views has been a critical task in computer vision and graphics. Recently, the trend of novel view synthesis is spearheaded by Neural Radiance Fields (NeRFs) [27] and its variants [1, 2], which achieves photo-realistic rendering quality. NeRF represents scenes with Multi-Layer Perceptron (MLP) and renders novel views using volumetric ray-marching. However, these MLP-based methods contain a relatively large number of layers, leading to lengthy training and rendering times, which still fall significantly short of real-time rendering. To accelerate rendering, many methods resort to auxiliary data structures, such as hash grids [28], points or voxel [39, 9, 38, 41]. However, due to the dense sampling of points along rays, they still face challenges in achieving real-time performance.
With a substantial demand for high-speed and photo-realistic rendering effects, 3D Gaussian Splatting (3D-GS) [17] has emerged as an efficient representation that can be rendered at high speed on the GPU. It has departed from the implicit scene representation using MLPs and instead opted for an explicit representation using Gaussian functions. As a result, 3D-GS avoids the expensive cost of densely sampling points in volume rendering, thus achieving real-time performance. Given a set of images with the camera parameters calibrated by SfM [33], 3D-GS aims to optimize a set of 3D Gaussians as graphical primitives to explicitly represent the scene. To render 2D images, these 3D Gaussians are projected onto the image plane ($z=1$ plane) via a projection function denoted as $\phi$ for differentiable rasterization.
@@
Unfortunately, Gaussian functions are closed under affine transformations but not under projective transformations. Therefore, 3D-GS adopts a local affine approximation [46], specifically approximating the projection function with the first two terms of its Taylor expansion. Nevertheless, approximations introduce errors, and the local affine approximation similarly contributes to these errors which may lead to artifacts in the rendered images. This paper exploits the relationship between the error of 3D-GS and the Gaussian mean through the analysis of the Taylor remainder term. Additionally, it identifies the circumstances under which the error is minimized by determining the extremum of the error function.
Finally, based on the extremum analysis of the error function, we propose a novel optimal projection method to minimize the projection error in 3D-GS. This method requires only minor code modifications, does not affect real-time rendering performance, and yet achieves a significant improvement in rendering quality. Specifically, we project along the direction from the Gaussian mean to the camera center, where the projection plane is tangent to the line connecting the Gaussian mean and the camera center. Compared with the straightforward approach of projecting 3D Gaussians onto the image plane of the camera, the proposed projection reduces expected error while maintaining real-time and realistic rendering performance. Furthermore, we find that the projection errors in 3D-GS tend to escalate with the continuous expansion of the field of view and the corresponding reduction in focal length. This increasing error leads to more and larger elongated Gaussians, as well as cloud-like Gaussian artifacts, consequently degrading the overall image quality. In contrast, our method exhibits greater realism and robustness. As illustrated in Figure 1, our method significantly reduces blurriness and artifacts caused by the projection error. Through experiments on various datasets and scenes, our method consistently outperforms the original projection method, particularly in settings with short focal lengths. And the method is easily implemented and adaptable to various camera models. To summarize, we provide the following contributions:
@@
  • We provide thorough error analysis for 3D-GS that may lead to artifacts and consequently degrade rendering quality. We have identified the correlation between this error and the Gaussian position.
@@
  • We propose a novel optimal projection capable of adapting to diverse camera models that adopts different tangent plane projections based on the directions from each Gaussian mean to the camera center instead of naively projecting all Gaussians onto the same plane
@@ 2 Related Work
Novel view synthesis is a longstanding challenge in computer vision and graphics. From traditional techniques  [33, 12, 20, 3, 35, 11, 6], to neural network-based scene representations [16, 31, 10, 26, 30, 34, 19], various approaches have struggled to address the problem of synthesizing a new view from captured images.
@@ 2.1 Neural Radiance Fields
The Neural Radiance Field (NeRF) [27] stands out as a successful neural rendering method based on MLPs, primarily owing to its encoding of position and direction. This encoding allows for effective reconstruction of high-frequency information in scenes. Notable improvements on this encoding have been made by MipNeRF [1], NeRF-W [24], FreeNeRF [43], and Instant NGP [28]. These enhancements enable the handling of multi-resolution image inputs, multi-illumination with occlusion image inputs, sparse-view inputs, and achieve nearly real-time rendering capabilities, respectively. Barron et al. introduced MipNeRF360 [2] as an extension of MipNeRF [1] to address the issue of generating low-quality renderings for unbounded scenes in NeRF. Importantly, these methods thoroughly exploit the intrinsic capabilities of NeRF as an implicit scene representation without introducing additional model priors.
However, due to the implicit representation of scenes and the dense sampling of points along rays, they still face challenges in achieving real-time performance.
@@
3D Gaussian Splatting (3D-GS) [17] has demonstrated notable advancements in rendering performance by departing from MLPs and ray sampling, opting instead for Gaussian functions and Gaussian Splatting. This paradigm shift has attracted considerable attention within the industry, leading to various studies building upon this methodology [7].
3D-GS finds application in SLAM systems [25], scene editing [8] and segmentation [4]. Efforts [21] have also been made to integrate 3D-GS with popular diffusion models [15, 36, 37] and dynamic scenes [23]. Recent endeavors aim to enhance the robustness in sparse-view scenarios [40, 42], performance, storage efficiency [29] and mesh reconstruction [13] of 3D-GS. However, these improvements do not specifically address errors associated with Gaussian projection. Potential methods for enhancing Gaussian functions include [44, 32]. Nevertheless, they introduce a Gaussian filter and a novel Gaussian depth sorting algorithm, respectively, without investigating the projection function and associated errors. These approaches still rely on projecting onto the $z=1$ plane.
This paper aims to explore the potential of 3D Gaussian Splatting by analyzing errors that may arise during the projection process. The analysis will delve into the factors contributing to these errors and propose methods to mitigate and reduce them.
@@ 3 Preliminaries
Similar to NeRF [27], the input to 3D-GS consists of a set of images, together with the corresponding cameras calibrated by SFM [33]. However, in contrast to NeRF, 3D-GS takes the sparse point cloud generated during the SFM calibration process as input. From these points, it constructs a set of 3D Gaussians as graphical primitives explicitly representing the scene. Each 3D Gaussian is characterized by its position (mean) $\boldsymbol{\mu}$, covariance matrix $\boldsymbol{\Sigma}$, and opacity $\alpha$, and carries anisotropic spherical harmonics (SH) representing directional appearance component (color) of the radiance field, following standard practice [28, 9]. Subsequently, these 3D Gaussians are projected onto the image plane ($z=1$ plane) via a projection function for differentiable rasterization. Next, we will provide a detailed explanation of the Gaussian projection process in 3D-GS.
@@
In the world space, we define a 3D Gaussian function by $G$, characterized by its mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. To project the Gaussian onto the image plane, the initial step in 3D-GS involves an affine transformation of this Gaussian function from the world space to the camera space via a viewing transformation matrix $\mathbf{W}$, yielding a transformed Gaussian function:
@@
$\begin{split}G\left(\mathbf{x}\right)=\exp\left\{-1/2\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)^{\top}\left(\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}\right)^{-1}\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)\right\}\text{.}\end{split}$ (1)
@@
We denote $G^{{}^{\prime}}$ as the 3D Gaussian in the camera space, with $\mathbf{x}^{{}^{\prime}}=\mathbf{W}\mathbf{x}$, $\boldsymbol{\mu}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}$ as the point, mean and covariance matrix in the camera space, respectively.
Subsequently, traditional 3D-GS requires projecting the Gaussian function in the camera space onto the $z=1$ plane, with the projection function $\varphi$ defined as:
$\varphi\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{\top}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x_{0}}^{\top}\mathbf{x_{0}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{\top}\mathbf{x}^{{}^{\prime}}\right)^{-1}$ (2)
@@
where $\mathbf{x_{0}}=\left[\begin{matrix}0,0,1\end{matrix}\right]^{\top}$ represents the projection of the camera space’s origin onto this plane. Expanding this projection function to the first order using Taylor series yields:
@@
$\begin{split}\varphi\left(\mathbf{x}^{{}^{\prime}}\right)&=\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)+R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\\ &\approx\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (3)
@@
where $\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ is the Jacobian of the affine approximation of the projective transformation, denoted as $\mathbf{J}$. Applying this local affine approximation by neglecting the Taylor remainder term allows to derive the 2D Gaussian function $G_{\text{2D}}$ projected onto the $z=1$ plane as:
@@
$\begin{split}G_{\text{2D}}\left(\mathbf{x}^{{}^{\prime}}\right)&=\exp\left\{-1/2\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)\right\}\\ &\approx\exp\left\{-1/2\left(\varphi\left(\mathbf{x}^{{}^{\prime}}\right)-\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\varphi\left(\mathbf{x}^{{}^{\prime}}\right)-\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right\}.\end{split}$ (4)
Since the rank of the matrix $\mathbf{J}$ is 2, the inverse of $\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ is, in fact, the inverse of the covariance matrix of the 2D Gaussian obtained by skipping the third row and column [46]. We denote $\mathbf{x}_{\text{2D}}=\varphi\left(\mathbf{x}^{{}^{\prime}}\right)$, $\boldsymbol{\mu}_{\text{2D}}=\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ and $\boldsymbol{\Sigma}_{\text{2D}}=\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ as the point, mean and $2\times 2$ covariance matrix of $G_{\text{2D}}$ in the image space, respectively. For a pixel $(u,v)$, we obtain the value of the 2D Gaussian function by querying $\mathbf{x_{\text{2D}}}=\left[\begin{matrix}u,v\end{matrix}\right]^{\top}$, and then perform alpha blending to derive the color of the pixel. The blue box in Figure 3 illustrates the projection process.